Terrell finds that if he takes a cruise in July it will cost $1200. This is 25% less than if he took the same cruise in December. What is the cost of taking the cruise in December?
Answer by oberobic(2304)

x = price in December
.75x = price in July
.75x = $1200
x = $1600
.
Check
(1200-1600)/1600 * 100 = -400/1600 * 100 = -.25*100 = -25%
Correct.
.
Done.